OpenStudy (anonymous):

The Pharaoh Chewing Gum Company has decided to sponsor an additional prize in the radio station's contest. They are giving away backstage passes for the concert! Pharaoh Chewing Gum manufactures a new product they are trying to promote. The new product is a pyramid-shaped gum with a square base. In the spirit of the other challenges, the company has decided to place their pyramid-shaped gum inside a clear glass giant bubble-gum shaped sphere. Each piece of gum has a base

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@wolf1728

Directrix (directrix):

Please post the remainder of the question. I think it cut off at "...of gum has a base."

OpenStudy (anonymous):

The Pharaoh Chewing Gum Company has decided to sponsor an additional prize in the radio station's contest. They are giving away backstage passes for the concert! Pharaoh Chewing Gum manufactures a new product they are trying to promote. The new product is a pyramid-shaped gum with a square base. In the spirit of the other challenges, the company has decided to place their pyramid-shaped gum inside a clear glass giant bubble-gum shaped sphere. Each piece of gum has a base measurement of 1 inch and a height of 0.75 inches. The glass sphere container has a diameter of 17.25 inches. How many pieces of Pharaoh Chewing Gum can fit inside the glass sphere? You must show all work to receive credit.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

This working of the problem is by Jordan Ph and is posted at: http://prezi.com/tvkjxwmmdlei/803-module-quiz/ Read the Penney Problem and Tennis Trouble. Then, decide what you think about this: Sphere: Diameter 17.25 V=4/3 (3.14*r³) 4/3 (3.14*8.625³) 4/3(3.14*641.619140625) 4/3 (2014.6841015625) Volume=2686.24546875 Gum: Base 1 in/ Height .75 V=1/3 (B) (h) 1/3 (1*1*1*1) (.75) 1/3 (.75) .25 Divide the sphere by the volume of the gum and get: 10,744 (approximately) pieces of gum @carlosortega